OpenStudy (anonymous):

There are 8 people on the debate team. In how many ways can the coach choose 4 members to send to competition? answers 70 1,680 40,320

OpenStudy (wolf1728):

The formula for combinations is: n! / k!*(n-k)! n = 8 k = 4 8! / 4!*4! (8*7*6*5*4*3*2*1) / 24 * 24 (8*7*6*5) / 24 =2 * 7 * 5 = 70

OpenStudy (anonymous):

Thank you !